OpenStudy (anonymous):

Write the equation of a line that passes through the points ( -2, 5) and (-2, 11).

OpenStudy (anonymous):

Are used to equations of the form \(f(x) = mx+b\) ? Or do you need a basic explanation how to do it?

OpenStudy (anonymous):

either or @cap_n_crunch

OpenStudy (anonymous):

Okay lets start with a quick way and if you have further quesions just ask :)

OpenStudy (anonymous):

the f(x)=mx=b would be fine i'm pretty sure @cap_n_crunch

OpenStudy (anonymous):

so first we need to get the slope \(m\) of the line. This can be done by simply dividing the offset in \(y\) by the offset in \(x\):\[ m = \frac{\Delta y}{\Delta x} = \frac{y_1 - y_2}{x1- x_2} = \frac{-2-(-2)}{5-11} = 0 \] As the slope is zero, \(x\) dissappears and we're left with a parallel line to the y-axis ( this can also be seen by the equal \(x\) coordinates of the two points ) Therefore the line is defined by:\[ y = -2 \]

OpenStudy (anonymous):

Thank you! @cap_n_crunch

OpenStudy (anonymous):

Your welcome :)